OpenStudy (anonymous):

If the mass of 0.25 mole of an element X is 2.25 g , the mass of one atom of X is

OpenStudy (anonymous):

@JFraser

OpenStudy (aaronq):

use these two formulas to interconvert from atoms, moles mass \(n=\dfrac{m}{M}\) and \(n=\dfrac{N}{N_a}\) n=moles m=mass M=molar mass N=number of particles/atoms/molecules \(N_A\)=avogadros number

OpenStudy (anonymous):

wait i am trying

OpenStudy (aaronq):

so you wanna find how many atoms there are (with the second formula), then divide the mass by the number of atoms.

OpenStudy (anonymous):

i got molar mass as 9

OpenStudy (aaronq):

well you don't really need the molar mass, i just posted that formula for completion. Follow what i said in my previous post.

OpenStudy (joannablackwelder):

I did this one like a conversion problem. 1 atom of X * (1 mole/6.02x10^23 atoms)*(2.25 g X/0.25 mole X) = ? g X